# Epoch-based Optimistic Concurrency Control in Geo-replicated Databases

### v1

@@ 5.1. Experimental Setup
We deploy the Minerva cluster on the Compute Canada Cloud. Each replica operates as a monolithic process, colocating storage and execution components within a single Virtual Machine (VM). Each node is provisioned with 4 vCPUs (Intel Xeon Skylake) and 48 GB of RAM. The operating system is Ubuntu 24.04, except for the GeoGauss baseline, which requires CentOS 7.6. All nodes interact via a high-speed LAN with 3 Gbps bandwidth. To prevent client-side bottlenecks, client threads are running on a separate 16-core VM.
@@
We compare Minerva against state-of-the-art systems that share key features: (1) Geo-replicated Multi-leader Databases, including GeoGauss (zhougeogauss2023), Ocean Vista (fanOceanVistaGossipbased2019), and CockroachDB (taftCockroachDBResilientGeoDistributed2020a); (2) Epoch-based Systems, including COCO (luEpochbasedCommitReplication2021); and (3) Deterministic Databases, including CalvinDB (thomsonCalvinFastDistributed2012; CalvinDB).
For GeoGauss, we configure the default $10ms$ epoch time as suggested in the paper. For COCO, we set the epoch time to the default $10ms$ and use COCO’s logical clock mode, which offers better performance. CalvinDB’s sequencer batch time is defaulted to $10ms$ as well. For a fair comparison, we configured all systems for full database replication by setting the number of replicas equal to the number of VM nodes, regardless of their individual sharding capabilities. Finally, all systems were set to in-memory mode, with disk persistency disabled when possible.
Since the source code for Ocean Vista is not publicly available, we reimplement its protocol using Minerva’s network and storage framework. While the original design relies on an epidemic gossip protocol for metadata dissemination, we simulate this behavior using a simplified periodic batch broadcast mechanism. Additionally, we replace its reliance on a synchronized clock using Network Time Protocol (NTP) with a logical deterministic ordering scheme derived from local clocks and replica IDs.
@@ 5.2. Benchmarks and Metrics
TPC-C is a sophisticated OLTP benchmark with more complex transactions that simulates the operation of a warehouse-based ordering system. We configure a 100-warehouse setup with $50\%$ of Payment transactions and $50\%$ of New-Order transactions. These are the two primary transaction types in the standard TPC-C benchmark, and this workload is used for evaluating all baselines except for CalvinDB. This workload is more computationally heavy and exhibits more contention than YCSB, but has smaller transaction sizes.
The main performance metrics are peak throughput (txns/s) and end-to-end latency (ms). Both metrics reflect only the successfully committed transactions at the client’s side. To determine peak throughput, we incrementally scale the client load until system saturation is observed. We then compute the average of highest sustainable throughput of five 30-second experiment runs. The specific client concurrency is tuned for each configuration to identify the maximum saturation point, ranging from 2,400 clients per replica in 3-node clusters to 1,600 clients per replica in 15-node clusters.
5.3. Effects of Cross-replica Latency
Figure 6. Throughput vs. Cross-replica Latency
We first evaluate the performance of the systems in geo-distributed scenarios by introducing synthetic network delay among 3 replicas using the Linux’s tc and netem tools. We vary the delay from $50ms$ to $200ms$, adding $5ms$ to $40ms$ of jitter to simulate various cross-region deployment conditions.
Figure 6 illustrates the throughput results. In a low-delay LAN environment, Minerva, CalvinDB, and GeoGauss establish the top performance level, each exceeding $10,000$ txns/s; notably, Minerva surpasses the latter two by approximately $30\%$. In contrast, COCO, CockroachDB, and Ocean Vista exhibit significantly lower throughput below $5,000$ txns/s, indicating architectural limitations under full replication. For instance, COCO is optimized for sharded execution, this configuration effectively treats every transaction as a cross-partition transaction, requiring coordination among all replicas and mitigating its performance optimizations.
Introducing $50$ ms of cross-replica latency degrades performance globally, yet underscores Minerva’s resilience. At this point, GeoGauss loses $33\%$ of its throughput and CalvinDB loses $40\%$, whereas Minerva’s throughput decreases by only about $15\%$, demonstrating the efficiency of its asynchronous replication and decoupled commit protocol. At $200ms$ delay, Minerva outperforms CalvinDB and GeoGauss by $3.8\times$.
With YCSB-A, Minerva’s advantage is less pronounced, where it only has a slight advantage over CalvinDB without network delay. However, as the delay increases, Minerva again shows superior performance, achieving $2.5\times$ the throughput of CockroachDB, the second-best baseline at $200ms$ latency.
@@ 5.3. Effects of Cross-replica Latency
GeoGauss and Ocean Vista exhibit lower median latencies than Minerva and CalvinDB, effectively achieving single-round-trip (1 RTT) commits whenever a replica is able to successfully receive commit information from all other replicas promptly. Transaction commit latency grows linearly with the increased cross-replica latency for all systems except for CockroachDB, which is severely affected by the network latency among replicas.
@@ 5.4. Scalability
We vary the number of replicas from 3 to 15 and measure the resulting throughput and latency for scalability.
@@
Notably, GeoGauss is the only system to experience throughput degradation when scaling from 3 to 15 replicas, whereas all others maintain or improve performance. This is likely attributed to its all-to-all communication pattern during the synchronization phase, which introduces overhead as the cluster expands.
Figure 9. Transaction Commit Latency vs. Scalability
@@
The latencies of individual transactions are shown in Figure 9. As expected, Minerva’s latency only increases slightly from $17ms$ to $18ms$ in TPC-C and from $18ms$ to $18.2ms$ in YCSB-A when scaling from 3 to 15 replicas. Most other baselines experience a steady latency except for COCO. Regarding Ocean Vista, we observe erratic fluctuations in tail latency (p95 and p99) despite a stable median, a behavior stemming from the inherent unpredictability of its dissemination protocol.
We see a similar trend for both scalability and network latency experiments between CalvinDB and Minerva. This is because both systems rely on deterministic execution. While Minerva’s OCC and optimized replication mechanism offered an edge in raw performance metrics, the trend remains similar, especially under high load, because Minerva behaves more like a deterministic database when contention is high due to the high-contention mode.
@@ 5.5. Contention Analysis
Figure 10 depicts the peak throughput under different Zipfian coefficients from $0.3$ to $1.7$ to change the contention of YCSB-A workload. This experiment also evaluates the performance trade-offs between the greedy approximate solver and the exact ILP solver within a zero-latency, 3-replica environment. Because CalvinDB does not support setting a Zipfian coefficient for contention, we simulate contention by setting a certain percentage of hot keys in the key space. COCO and CockroachDB  [^4] CockroachDB only supports single operation YCSB benchmark; the 10-operation YCSB transactions are emulated using the kv workload in other experiments, which does not support varying contentions (cdbworkload). are excluded from this experiment because they do not provide configurable contention control.
@@
We observe about 5% of all transactions are re-executed when the coefficient is set 0.3, 30% when the coefficient is 0.7, and maxed out at about 35% for higher contention because the high-contention mode is triggered and preventing more transactions being optimistically executed then re-executed.
The baseline CalvinDB does not experience any performance change as expected, since the deterministic execution model is largely unaffected by contention levels. GeoGauss and Ocean Vista both experience a performance drop as contention increases, contributing to a higher abort rate in their optimistic execution model.
@@ 5.6. Fault Tolerance
We evaluate Minerva’s resilience to replica failures by simulating crash failures in 5-replica clusters running the TPC-C benchmark. We introduce failures to the coordinator or a random replica at the 15-second mark during a 30-second experiment run by shutting down the server processes. The clients are not connected to the failed replica and the overall system load is set to about 50% of the peak throughput with 400 clients across the remaining replicas to avoid saturation and better observe the recovery process.
As shown in Figure 11, when the coordinator fails, the commit throughput is temporarily disrupted for about 1 seconds, which correspond to the Raft leader election in our implementation ($400ms$ heartbeat). After the new leader is elected, the system resumes normal operation without any performance degradation. When a random replica fails, the system continues to operate with a brief drop in throughput. This is because the Raft protocol in our implementation (DotNext) waits for a timeout ($50ms$) for all replica to acknowledge, then it checks for quorum instead of immediately committing as soon as a majority vote is received. However, once the system detects the failure, it removes the failed replica from the cluster configuration, and the throughput recovers to normal levels without further degradation. These results demonstrate Minerva’s ability to maintain high availability and performance in the face of replica failures, leveraging its underlying Raft-based consensus mechanism for fault tolerance.
@@ 5.7. Performance breakdown
We profile Minerva’s commit process with two experiment runs using the YCSB-A workload with 3 replicas and no cross-replica latency; the system is saturated with average end-to-end transaction latency at about $500ms$ (at this point, the system throughput no longer increases with more clients). We measure a high contention workload (about $48\%$ conflicting and stale operations) and a low contention workload (about $2\%$ conflicting and stale operations). We disable the high-contention mode so that all transactions go through the full commit process for more accurate analysis of the commit process.
@@
Table 1 presents the average operation time breakdown for the serial commit process under both low-contention and high-contention scenarios. In both scenarios, the total commit time remains relatively consistent at approximately $32ms$. However, the distribution of time across different operations varies significantly between the two scenarios.
In the low-contention scenario, the majority of the commit time is consumed by applying the OCC results (i.e., commit the write-sets of transactions). This is expected, as most transactions pass local OCC validation and are committed directly. Conversely, in the high-contention scenario, the dominant factor in commit time is the deterministic re-execution of transactions. This reflects the high rate of conflicts and subsequent aborts, leading to a substantial portion of the workload being re-executed.
We also observed that time-consuming operations consistently involve accessing the storage layer, even in our in-memory configuration which only requires memory copying. For instance, checking for stale transactions requires reading the latest database snapshot version for every key in the read-set, and conflict graph construction involves building complex in-memory data structures based on all read and write keys. Surprisingly, the MWIS solving phase consumed a relatively steady, small amount of time across both scenarios, even when using the exact solver. This shows that the overhead is actually memory-bound in data access rather than computation-bound by our algorithms.
@@ 5.8. Epoch Time
Figure 12. Epoch Time, Network Latency and batch Generate Timeout
Figure 12 illustrates the impact of varying epoch times on system throughput under different network latencies and batch generation times for TPC-C workload with 3 replicas. The batch generation timeout is the maximum time a replica waits before broadcasting a batch, even if the batch size threshold has not been met.
Despite that we expected the optimal performance should be gained by matching the epoch time with the network latency, the results indicate that shorter epoch times generally yield better throughput, regardless of network latency. This is because epoch time only determines the generation frequency of the global commit proposals, while the actual commit is not synchronized to the epoch boundaries. Thus, shorter epoch times allow for more frequent commit opportunities, leading to improved throughput. Similarly, shorter batch generation time also leads to better performance, as it reduces the waiting time for transactions to be included in a batch, thereby improving overall system throughput.
5.9. Observation and Discussion
We observe that Minerva’s algorithm is bottlenecked by the serialized commit process. Even though the commit process requires no coordination, and it is asynchronous to the execution and replication. It is still on the critical path for end-to-end transaction life cycle and thus established the lower bound of latency. This is essentially a producer-consumer problem where the system is limited by the rate at which the serialized commit handler can consume the optimistically produced transactions results. Therefore, regardless of how fast the local OCC execution is, the overall throughput cannot exceed the commit throughput.
@@ 5.9. Observation and Discussion
A potential improvement is to support weaker isolation and consistency levels, such as in RedBlue Consistency (li_making_2012) or Janus CRDT (maoMakingCRDTsNot2024a). With batch-committed transactions, some conflicting operations could be resolved through commutative operations, such as those provided by Conflict-Free Replicated Data Types (CRDTs) (shapiroConflictFreeReplicatedData2011), or by using Last-Writer-Wins (LWW) conflict resolution, rather than resorting to re-execution. Weakly consistent write also enables a fast-path commit, where transactions can be committed as soon as they gain a PoA because it is guaranteed to be persisted already.
@@
Another drawback of replicating both the transaction input and the write-sets is the resulting increased bandwidth consumption. This raises cost concerns and could potentially degrade performance if the network connection bandwidth is limited. Compression techniques could be employed to mitigate this issue.
Finally, like Calvin and other deterministic databases, Minerva requires transactions to be predefined with known read and write sets and thus only supports one-shot transactions for deterministic re-execution. However, we could utilize the OCC executor for interactive transactions by using only the OCC execution path with the same replication and conflict detection mechanism, albeit with a possibility for transactions to be aborted.
@@ 6. Related Work
Calvin (thomsonCalvinFastDistributed2012) is a well-known deterministic database that first introduces the deterministic concurrency control. Calvin employs a distributed sequencer layer where each database shard is assigned a sequencer to stamp arriving transactions with a global sequence number. Sequencers collect transactions into batches independently, and the commit phase is based on network-synchronized clock. Finally, transactions in the batch are executed deterministically according to their sequence number. To maximize throughput, concurrent data access during execution is managed by a deterministic locking protocol.
While Calvin does not support multi-leader transaction execution, every replica can accept client requests. If synchronous replication is enabled, these requests are replicated to the sequencers via a Paxos-based consensus protocol (implemented with ZooKeeper) to ensure durability before execution.
Aria (luAriaFastPractical2020) is a recent distributed deterministic database that utilizes snapshots for optimistic execution on batches of transactions. This design facilitates high concurrency without requiring a priori knowledge of read/write sets, in contrast to systems relying on deterministic locking or dependency analysis. Aria also has an execution phase and a validation phase, which identifies and aborts conflicting transactions. However, aborted transactions must be rescheduled in a subsequent batch. In high-contention scenarios, this leads to performance degradation and necessitates a synchronous re-execution fallback, a strategy similar to the conflict resolution mechanism in Minerva. Like Calvin, Aria does not inherently support multi-leader replication.
The main distinction between Minerva and these deterministic systems is that Minerva replicates both transaction inputs and execution results (write-sets). This dual-replication approach enables faster commit time for non-conflicting transactions.
@@
GeoGauss (zhougeogauss2023) is a recent geo-replicated OLTP database built on openGauss (openGauss) that supports multi-leader replication. Like Minerva, GeoGauss is OCC based and replicas exchange their write-sets with all peers at the end of each epoch. These sets are then merged into a final commit set using deterministic, commutative merge rules derived from delta-state CRDTs (almeida_delta_2018). GeoGauss resolves conflicts by deterministically aborting one of the contending transactions based on sequence numbers. Minerva’s MWIS-based conflict resolution optimizes this process by reducing the number of aborted/re-executed transactions.
Additionally, GeoGauss’s epoch boundary functions as a global synchronization barrier, requiring every replica to await updates from all other replicas. While it utilizes Raft for configuration management to monitor replicas, such global synchronization may cause uncertain blocking delays upon failures.
@@
Hackwrench (dongFineGrainedReExecutionEfficient2023) uses a two-tiered, batch-based commitment protocol. For conflict resolution, Hackwrench re-executes aborted transactions based on a predefined dataflow/dependency graph, contrasting with Minerva’s runtime dependency analysis. Hackwrench’s replication is achieved through a leaderless quorum-based write to storage nodes, effectively decoupling the transaction execution plane from the storage replicas. Neither COCO nor Hackwrench considers multi-leader replication.
@@
Ocean Vista (OV) enables multi-leader geo-replication through a unified visibility control protocol (fanOceanVistaGossipbased2019). It decouples replication from commitment with a two-phase approach: an initial S-phase uses lightweight consensus to replicate transaction ”functors” (logic and inputs) across a quorum, similar to deterministic databases’ transaction input replication. Final execution and visibility (akin to actually committing the transactions so clients can see the result) decisions occur in an asynchronous E-phase, where nodes independently apply transactions only after a global gossip protocol establishes a stability watermark, ensuring all prior transactions are replicated.
The order of transactions is determined by timestamps synchronized across replicas using NTP, so OV optimizes for visibility latency but is sensitive to clock skew; it still requires consensus to ensure that transactions are replicated properly.
@@
Epidemic algorithms (holliday2003epidemic) use only asynchronous gossip protocols to propagate updates and maintain consistency without a centralized coordinator to maximize availability. They rely on causal ordering and local stability verifications to ensure serializability. However, this approach may experience unpredictable convergence times due to the probabilistic nature of message dissemination. Furthermore, ensuring strict consistency requires nodes to await updates from all peers to detect concurrent conflicts, where a single slow node or network failure can indefinitely stall progression.
Other Unified Protocols
@@
TAPIR (zhang_tapir_2018) abandons linearizable replication in favor of an inconsistent replication protocol with an OCC-style validation, though it still relies on client execution and client-side 2PC for atomic commitment. The C&C framework (maiyyaCNC2019) combines atomic commitment and replication by piggybacking commit decisions onto consensus messages, yet it still requires a separate distributed concurrency control mechanism. Similarly, Janus (mu_consolidating_2016) combines concurrency control and replication by tracking dependency graphs to detect conflicts, invoking consensus only when cycles exist. However, none of these systems achieve the complete unification of concurrency control, replication, and atomic commitment found in Minerva. Furthermore, they often target different communication models or fault-tolerance guarantees and lack native support for the multi-leader architecture.
7. Conclusion

### v2

@@ 5.1. Experimental Setup
We deploy the Minerva cluster on the Compute Canada Cloud. Each replica operates as a monolithic process, colocating storage and execution components within a single Virtual Machine (VM). Each node is provisioned with 4 vCPUs (Intel Xeon Skylake) and 48 GB of RAM. The operating system is Ubuntu 24.04, except for the GeoGauss baseline, which requires CentOS 7.6. All nodes interact via a high-speed LAN with 3 Gbps bandwidth. To prevent client-side bottlenecks, client threads run on a separate 16-core VM.
@@
We compare Minerva against state-of-the-art systems that share key features: (1) Geo-replicated databases, including GeoGauss (zhougeogauss2023), Ocean Vista (fanOceanVistaGossipbased2019), and CockroachDB (taftCockroachDBResilientGeoDistributed2020a); (2) epoch-based systems, including COCO (luEpochbasedCommitReplication2021); and (3) deterministic databases, including CalvinDB (thomsonCalvinFastDistributed2012; CalvinDB).
For GeoGauss, COCO, and Calvin, we configure all to use the default $10ms$ epoch time as suggested in their papers. We also use COCO’s logical clock mode, which offers better performance. For a fair comparison, we set all systems for full database replication by setting the number of replicas equal to the number of VM nodes, regardless of their individual sharding capabilities. Finally, all systems are set to in-memory mode, with disk persistency disabled when possible.
The source code for Ocean Vista is not publicly available, so we reimplement its protocol using Minerva’s network and storage layer. While the original design relies on an epidemic gossip protocol for metadata dissemination, we simulate this behavior using a simplified periodic batch broadcast mechanism. Additionally, we replace its reliance on a synchronized clock using Network Time Protocol (NTP) with a logical deterministic ordering scheme derived from local clocks and replica IDs.
@@ 5.2. Benchmarks and Metrics
TPC-C is a sophisticated OLTP benchmark with more complex transactions that simulates the operation of a warehouse-based ordering system. We configure a 100-warehouse setup with $50\%$ of Payment transactions and $50\%$ of New-Order transactions. These are the two primary transaction types in the standard TPC-C benchmark, and this workload is used for evaluating all baselines except for CalvinDB. It is more computationally heavy and exhibits more contention than YCSB, but has smaller payload sizes.
The main performance metrics are peak throughput (txns/s) and end-to-end latency (ms). Both metrics reflect only the successfully committed transactions on the client side. To determine peak throughput, we incrementally scale the client load until system saturation is observed. We then compute the average of the highest sustainable throughput across five 30-second experiment runs. The specific client concurrency is tuned for each configuration to identify the maximum saturation point, ranging from 2000 to 8000 total client threads.
5.3. Effects of Cross-replica Network Latency
Figure 6. Throughput vs. Cross-replica Network Latency
We first evaluate the performance of the systems in geo-distributed scenarios by introducing synthetic network delay among 3 replicas using Linux’s tc and netem tools. We vary the delay from $50ms$ to $200ms$, adding $5ms$ to $40ms$ of jitter to simulate various cross-region deployment conditions.
Throughput
Figure 6 illustrates the throughput results. In a low-delay LAN environment, Minerva, CalvinDB, and GeoGauss achieve the highest performance levels, each exceeding $10,000$ txns/s; notably, Minerva surpasses the latter two by approximately $30\%$. In contrast, COCO, CockroachDB, and Ocean Vista exhibit significantly lower throughput, below $5,000$ txns/s, indicating architectural limitations under full replication. For instance, COCO is optimized for sharded execution; this configuration effectively treats every transaction as a cross-partition transaction, requiring 2PC among all replicas and limiting its performance optimizations.
Introducing $50$ ms of cross-replica latency degrades performance globally. At this point, GeoGauss loses $33\%$ of its throughput and CalvinDB loses $40\%$, whereas Minerva’s throughput decreases by only about $15\%$, demonstrating the efficiency of its asynchronous replication and decoupled commit protocol. At $200ms$ delay, Minerva outperforms CalvinDB and GeoGauss by $3.8\times$.
With YCSB-A, Minerva only has a slight advantage over CalvinDB without network delay. However, as the delay increases, Minerva again shows superior performance, achieving $2.5\times$ the throughput of CockroachDB, the second-best baseline at $200ms$ latency.
@@ 5.3. Effects of Cross-replica Network Latency
GeoGauss and Ocean Vista exhibit lower median latencies than Minerva and CalvinDB, effectively achieving single-round-trip commits whenever a replica is able to promptly receive commit information from all other replicas. Transaction commit latency grows linearly with increasing cross-replica latency for all systems except CockroachDB, which is severely affected by network latency among replicas.
@@ 5.4. Scalability
We vary the number of replicas from 3 to 15 and measure the resulting throughput and latency to evaluate scalability.
@@
Notably, GeoGauss is the only system to experience throughput degradation when scaling from 3 to 15 replicas, whereas all others maintain or improve performance. This is likely due to its all-to-all communication in the synchronization phase, which introduces overhead as the cluster expands.
Figure 9. Scalability in Transaction Commit Latency
@@
The latencies of individual transactions are shown in Figure 9. As expected, Minerva’s latency only increases slightly from $17ms$ to $18ms$ in TPC-C and from $18ms$ to $18.2ms$ in YCSB-A when scaling from 3 to 15 replicas. Most other baselines experience a steady latency except for COCO. Regarding Ocean Vista, we observe erratic fluctuations in tail latency (p95 and p99) despite a stable median, which may stem from the unpredictability of its dissemination protocol.
We see a similar trend in both the scalability and network-latency experiments for CalvinDB and Minerva. This is because both systems rely on deterministic execution. While Minerva’s OCC and optimized replication mechanism offer an edge in raw performance metrics, the trend remains similar, especially under high load, because Minerva behaves more like a deterministic database when contention is high due to the high-contention mode.
@@ 5.5. Contention Analysis
Figure 10 depicts the peak throughput under different Zipfian coefficients from $0.3$ to $1.7$ to change the contention of YCSB-A workload. This experiment also evaluates the performance trade-offs between the greedy approximate solver and the exact ILP solver in a LAN environment with $3$ replicas. Because CalvinDB does not support setting a Zipfian coefficient for contention, we simulate contention by setting a certain percentage of hot keys in the key space. COCO and CockroachDB  [^4] CockroachDB only supports single operation YCSB benchmark; the 10-operation YCSB transactions are emulated using the kv workload in other experiments, which does not support varying contentions (cdbworkload). are excluded from this experiment because they do not provide configurable contention control.
@@
We observe about 5% of all transactions are re-executed when the coefficient is set to 0.3, 30% when the coefficient is 0.7, and maxes out at about 35% for higher contention because the high-contention mode is triggered, preventing more transactions from being executed twice.
The baseline CalvinDB does not experience any performance change, as expected, since the deterministic execution model is largely unaffected by contention levels. GeoGauss and Ocean Vista both experience a performance drop as contention increases, contributing to a higher abort rate in their optimistic execution models.
@@ 5.6. Fault Tolerance
We evaluate Minerva’s resilience to replica failures by simulating crash failures in a 5-replica cluster running the TPC-C benchmark. We introduce failures to the coordinator or a random replica at the 15-second mark during a 30-second experiment run by shutting down the server processes. The clients are not connected to the failed replica and the overall system load is set to about 50% of the peak throughput with 400 clients across the remaining replicas to avoid saturation and better observe the recovery process.
As shown in Figure 11, when the coordinator fails, the commit throughput is temporarily disrupted for about 1 second, which corresponds to the Raft leader election in our implementation ($400ms$ heartbeat). After the new leader is elected, the system resumes normal operation without further performance degradation. When a random replica fails, the system continues to operate with a brief drop in throughput. This is because the Raft protocol in our implementation (DotNext) waits for a timeout ($50ms$) for all replicas to acknowledge, then checks for a quorum instead of immediately committing as soon as a majority vote is received. However, once the system detects the failure, it removes the failed replica from the cluster configuration, and the throughput recovers to normal levels. These results demonstrate Minerva’s ability to maintain high availability and performance in the face of replica failures, leveraging its underlying Raft-based consensus mechanism for fault tolerance.
@@ 5.7. Performance Breakdown
We profile Minerva’s commit process with two experiment runs using the YCSB-A workload with 3 replicas, no added latency, and high-contention mode disabled. The system is saturated with an average end-to-end transaction latency of about $500ms$ (at this point, the system throughput no longer increases with more clients). We measure a high-contention workload (about $48\%$ conflicting and stale operations) and a low-contention workload (about $2\%$ conflicting and stale operations).
@@
Table 1 presents the average operation-time breakdown for the serial commit process under both low-contention and high-contention scenarios. In both scenarios, the total commit time remains relatively consistent at approximately $32$ ms. However, the distribution of time across different operations varies significantly between the two scenarios.
In the low-contention scenario, the majority of the commit time is consumed by applying the OCC results (i.e., moving the write-sets to storage). This is expected, as most transactions pass the validation phase and are committed directly. Conversely, in the high-contention scenario, the dominant factor in commit time is the deterministic re-execution of transactions. This reflects the high rate of conflicts and subsequent aborts, leading to a substantial portion of the workload being re-executed.
We also observe that time-consuming operations consistently involve accessing the storage layer, even in our in-memory configuration, which only requires memory copying. For instance, checking for stale transactions requires reading the latest database snapshot version for every key in the read-set, and conflict graph construction involves building complex in-memory data structures based on all read and write keys. Surprisingly, the MWIS solving phase consumes a relatively steady, small amount of time across both scenarios, even when using the exact solver. This shows that the overhead is actually memory-bound because of data movement rather than computation-bound by our algorithms.
@@ 5.8. Epoch Time
Figure 12. Effect of Epoch Times and Batch Generation Timeouts with Varying Network Latencies
Figure 12 illustrates the impact of varying epoch times on system throughput under different network latencies and batch generation times for the TPC-C workload with 3 replicas. The batch generation timeout is the maximum time a replica waits before broadcasting a batch, even if the batch size threshold has not been met.
Although we expect optimal performance to be gained by matching the epoch time to the network latency, the results indicate that shorter epoch times generally yield better throughput, regardless of network latency. This is because the epoch time only determines the generation frequency of the global commit proposals, while the actual commit is not synchronized to the epoch creation time. Thus, shorter epoch times allow for more frequent commit opportunities, leading to improved throughput. Similarly, a shorter batch generation time also leads to better performance, as it reduces the waiting time for transactions to be included in a batch, thereby improving overall system throughput.
5.9. Observations and Discussion
We observe that Minerva is bottlenecked by the serialized commit process. Even though the commit process requires no coordination, and is asynchronous to execution and replication, it is still on the critical path for the end-to-end transaction life cycle and thus establishes a lower bound for latency. This is essentially a producer-consumer problem where the system is limited by the rate at which the serialized commit handler can consume the optimistically produced transaction results. Therefore, regardless of how fast the local OCC execution is, the overall throughput cannot exceed the commit throughput.
@@ 5.9. Observations and Discussion
A potential improvement is to support weaker isolation and consistency levels, such as in RedBlue Consistency (li_making_2012) or Janus CRDT (maoMakingCRDTsNot2024a). With batch-committed transactions, some conflicting operations can be resolved through commutative operations, such as those provided by Conflict-Free Replicated Data Types (CRDTs) (shapiroConflictFreeReplicatedData2011), or by using Last-Writer-Wins (LWW) conflict resolution, rather than resorting to re-execution. Weakly consistent writes also enable a fast-path commit, where transactions can be committed as soon as they gain a PoA because the transaction is guaranteed to be persisted with the PoA.
@@
Another drawback of replicating both input and results is the increased bandwidth consumption. This raises cost concerns and could potentially degrade performance if the network connection bandwidth is limited. Compression techniques could be employed to mitigate this issue.
Finally, like Calvin and other deterministic databases, Minerva requires transactions to be predefined with known read and write sets, and thus only supports one-shot transactions for deterministic re-execution. However, we can utilize the OCC executor for interactive transactions by using only the OCC execution path with the same replication and conflict detection mechanism, albeit with the possibility that transactions may be aborted.
@@ 6. Related Work
Calvin (thomsonCalvinFastDistributed2012) first introduces deterministic concurrency control. Calvin employs a distributed sequencer layer where each database shard is assigned a sequencer to stamp arriving transactions with a global sequence number. Sequencers collect transactions into batches independently, and the commit phase is based on a network-synchronized clock. Finally, transactions in the batch are executed deterministically according to their sequence number. To maximize throughput, concurrent data access during execution is managed by a deterministic locking protocol. While Calvin is not built for multi-leader transaction execution, every replica can still accept client requests and route them to the leader replica for replication via a Paxos-based consensus protocol (implemented with ZooKeeper) to ensure durability before execution.
Aria (luAriaFastPractical2020) proposes a deterministic execution algorithm that utilizes snapshots for optimistic execution on batches of transactions. This design facilitates high concurrency without requiring a priori knowledge of read/write sets, in contrast to systems relying on deterministic locking or dependency analysis. Aria also has an execution phase and a validation phase, which identifies and aborts conflicting transactions. However, aborted transactions must be rescheduled in a subsequent batch. In high-contention scenarios, this leads to performance degradation and necessitates a synchronous re-execution fallback, a strategy similar to our conflict resolution mechanism. Replication is not part of Aria’s protocol as in Minerva.
HDCC (HongHDCC2025) is a recent work that also combines OCC and deterministic concurrency control by rescheduling aborted transactions with a deterministic scheduler through known read/write sets. Like Aria, HDCC does not consider replication either. Caracal (qin_caracal_2021) also introduces a similar epoch-batch-based deterministic concurrency control protocol.
The main distinction between Minerva and these deterministic systems is that Minerva replicates both transaction inputs and the speculative execution results (write-sets). This dual-replication approach not only enables faster commit time for non-conflicting transactions, but also serves as a method for obtaining write-sets in deterministic execution, which otherwise requires pre-declaration or reconnaissance operations.
@@
GeoGauss (zhougeogauss2023) is a recent geo-replicated OLTP database built on openGauss (openGauss) that supports multi-leader replication. Like Minerva, GeoGauss is OCC-based, and replicas exchange their write-sets with all peers at the end of each epoch. These sets are then merged into a final commit set using deterministic, commutative merge rules derived from delta-state CRDTs (almeida_delta_2018). GeoGauss resolves conflicts by deterministically aborting one of the contending transactions based on sequence numbers. Minerva’s MWIS-based conflict resolution optimizes this process by reducing the number of aborted/re-executed transactions.
Additionally, GeoGauss’s epoch boundary functions as a global synchronization barrier, requiring every replica to await updates from all other replicas. While it utilizes Raft for configuration management to monitor replicas, such synchronization may cause uncertain blocking delays upon failures.
@@
Hackwrench (dongFineGrainedReExecutionEfficient2023) uses a two-tiered, batch-based commitment protocol. For conflict resolution, Hackwrench re-executes aborted transactions based on a predefined dataflow/dependency graph, contrasting with Minerva’s runtime dependency analysis. Hackwrench’s replication is achieved through a leaderless quorum-based write to storage nodes, effectively decoupling the transaction execution plane from storage. Neither COCO nor Hackwrench consider multi-leader replication.
@@
Ocean Vista (OV) enables multi-leader geo-replication through a unified visibility control protocol (fanOceanVistaGossipbased2019). It decouples replication from commitment with a two-phase approach: an initial S-phase uses lightweight consensus to replicate transaction ”functors” (logic and inputs) across a quorum, similar to deterministic databases’ transaction input replication. Final execution and visibility (akin to actually committing the transactions so clients can see the result) decisions occur in an asynchronous E-phase, where nodes independently apply transactions only after a global gossip protocol establishes a stability watermark, ensuring all prior transactions are replicated. The order of transactions is determined by timestamps synchronized across replicas using NTP, so OV optimizes for visibility latency but is sensitive to clock skew; it still requires consensus to ensure that transactions are replicated properly.
@@
Epidemic algorithms (holliday2003epidemic) use only asynchronous gossip protocols to propagate updates and maintain consistency without a centralized coordinator to maximize availability. They rely on causal ordering and local stability verifications to ensure serializability. However, this approach may experience unpredictable convergence times due to the probabilistic nature of message dissemination. Furthermore, ensuring strict consistency requires nodes to await updates from all peers to detect concurrent conflicts, meaning a single slow node or network failure can indefinitely stall progression.
Mako (shenMakoSpeculativeDistributed2025) uses a similar speculative execution model and asynchronous replication. It relies on primary-backup replication, but utilizes a deterministic replay model on the replicas to reach consistency rather than replicating the results. The 2PC is optimized by committing speculatively as long as the leader shards agree to commit.
Other Optimized Protocols
@@
TAPIR (zhang_tapir_2018) abandons linearizable replication in favor of an inconsistent replication protocol with an OCC-style validation, though it still relies on client execution and client-side 2PC for atomic commitment. The C&C framework (maiyyaCNC2019) combines atomic commitment and replication by piggybacking commit decisions onto consensus messages, yet it still requires a separate distributed concurrency control mechanism. Similarly, Janus (mu_consolidating_2016) combines concurrency control and replication by tracking dependency graphs to detect conflicts, invoking consensus only when cycles exist.
Compared to Minerva, these systems target different communication models or fault-tolerance guarantees, or a specific layer, rather than a unified protocol for the multi-leader architecture.
7. Conclusions